OpenStudy (lalaly):

Find the absolute extreme values of f(x, y) = x^2 +x+y^2 −2y defined on { x^2 + y^2 

OpenStudy (anonymous):

not the best way of doing it

OpenStudy (anonymous):

i think u left something out

OpenStudy (anonymous):

you set grad f =0 for inside the disc

OpenStudy (anonymous):

grad f = ( 2x +1 , 2y-2) =zero vector so (-1/2 , 1 ) is a candidate for max /min

OpenStudy (anonymous):

then let g = x^2 +y^2 (the boundary ) by lagrange eqn ( 2x +1 , 2y-2 ) = A ( 2x, 2y ) so 2x +1 = 2Ax -- eqn 1 2y-2 = 2Ay --- eqn 2 now solve both eqns for A, and equate (2x+1) / (2x) = (2y-2)/ (2y) (2x+1)/(2x) = (y-1)/ y y(2x+1) = 2x(y-1) so y=-2x sub this into the boundary of the constraint x^2 +4x^2 = 1 x= +- sqrt(1/5)

OpenStudy (anonymous):

so then you have two more points ( sqrt(1/5) , - 2 sqrt(1/5) ) and ( -sqrt(1/5) , 2sqrt(1/5) )

OpenStudy (anonymous):

so you have three points , sub them into the function to find the respective function values , then chose the largest and smallest values

OpenStudy (anonymous):

note: the point (-1/2, 1 ) is outside the unit disc , so discard that .

OpenStudy (blacksteel):

Yeah, elec is right - I tried to do this problem an easier way but skipped a step which gives min/max points that aren't extrema of x and y.

OpenStudy (lalaly):

oohh thankyou elecengineer.....that helpd alot and thankyouu blacksteel :D